# SAGI: Semantically Aligned and Uncertainty Guided AI Image Inpainting

### v2

@@ Abstract
Recent advancements in generative AI have made text-guided image inpainting—adding, removing, or altering image regions using textual prompts—widely accessible. However, generating semantically correct photorealistic imagery, typically requires carefully-crafted prompts and iterative refinement by evaluating the realism of the generated content - tasks commonly performed by humans. To automate the generative process, we propose Semantically Aligned and Uncertainty Guided AI Image Inpainting (SAGI), a model-agnostic pipeline, to sample prompts from a distribution that closely aligns with human perception and to evaluate the generated content and discard one that deviates from such a distribution, which we approximate using pretrained Large Language Models and Vision-Language Models. By applying this pipeline on multiple state-of-the-art inpainting models, we create the SAGI Dataset (SAGI-D), currently the largest and most diverse dataset of AI-generated inpaintings, comprising over 95k inpainted images and a human-evaluated subset. Our experiments show that semantic alignment significantly improves image quality and aesthetics, while uncertainty guidance effectively identifies realistic manipulations — human ability to distinguish inpainted images from real ones drops from 74% to 35% in terms of accuracy, after applying our pipeline. Moreover, using SAGI-D for training several image forensic approaches increases in-domain detection performance on average by 37.4% and out-of-domain generalization by 26.1% in terms of IoU, also demonstrating its utility in countering malicious exploitation of generative AI. Code and dataset are available at https://github.com/mever-team/SAGI
@@ 1 Introduction
Figure 1: The first row illustrates SAOR: an input image is processed to identify objects (via labels or segmentation) and generate a caption. The language model selects an object and generates a prompt, which, along with the image and mask, guides the inpainting model to produce the result. The second row depicts UGDA: if the inpainted image passes an initial realism check, it undergoes a second evaluation. The VLM compares the inpainted and original images twice, in reverse order. If the VLM’s responses differ (indicating uncertainty) or both favor the inpainted image, it is labeled as deceiving; otherwise, undeceiving.
@@
Image inpainting—reconstructing regions in images—has become highly accessible with powerful generative AI tools, enabling non-experts to create photorealistic edits [84]. Text-guided inpainting, which adds, removes, or alters regions using textual prompts, has advanced significantly through models like Stable Diffusion [58], DALL-E [56], and Imagen [61]. These models enable edits using textual descriptions and masks. Creating high-quality results, however, often requires multiple attempts and iterative refinement [42], as even advanced models struggle to follow complex prompts, limiting their ability to consistently produce high-quality results [10]. However, successful generations can be so realistic that humans struggle to distinguish them from genuine photographs [10]. The realism of such AI-generated images, while impressive from a technical standpoint, also presents significant challenges regarding their potential misuse for malicious purposes [73, 78].
@@
Automated inpainting generation faces two key challenges. First, while state-of-the-art models excel at manipulation, they require detailed prompts for realistic results—basic object labels often lack sufficient context [41]. Second, automating realism assessment is difficult, as standard quality metrics may not reflect perceived realism. To address these, we propose SAGI (Semantically Aligned and Uncertainty Guided AI Image Inpainting), a unified framework that integrates Semantically Aligned Object Replacement (SAOR) and Uncertainty Guided Deceptiveness Assessment (UGDA). SAOR is a method that leverages the semantic content of an image and Large Language Models (LLMs) to create detailed, context-aware prompts that guide the inpainting process, producing results of higher aesthetic quality compared to simple object labels. UGDA is a realism assessment method, that uses vision-language models to compare inpainted images with their originals, helping identify convincing manipulations. We validate this approach through a user study, showing it aligns with human perception of image realism. The framework we propose enables a wide range of applications. First, it facilitates the generation of images with enhanced aesthetic quality and realism through an automated pipeline. Second, it enables the creation of forgery datasets that challenge existing forensic methods, identifying weaknesses as inpainting evolves and supporting the development of robust tools for detecting manipulations. An overview of the approach is presented in Fig. 1.
@@
Using this framework along with multiple state-of-the-art inpainting models, we create the SAGI-D (SAGI-Dataset) dataset for detecting AI-generated inpainting. The dataset contains 95,839 inpainted images generated from 78,684 original images from three datasets: MS-COCO [38], RAISE [18], and OpenImages [33]. We provide each inpainted image with its original version, inpainting mask, and text prompt. For thorough model evaluation, we create both in-domain and out-of-domain testing splits. The in-domain split uses the same LLM and source images as the training set, allowing assessment on familiar data. The out-of-domain split uses different source images and a different LLM, testing how well models handle new types of data.
@@
  • We introduce semantically aligned object replacement to generate semantically correct prompts.
@@
  • We present the SAGI-D dataset, the largest and most diverse dataset of AI-generated inpainted images to date.
@@ 2.1 Image Inpainting
Early image inpainting methods used diffusion for image restoration [8, 9, 7, 11]. Later methods adopted exemplar-based approaches for removing large objects [17, 24, 28, 21]. Recent methods are deep learning-based, using CNNs [6], autoencoders [31], Fourier Convolutions [65], Transformers [72], or Diffusion Models [13]. GAN-based methods have also been widely adopted, combining autoencoders and GANs to generate coherent images, with techniques such as Context Encoders [53] and dilated convolutions improving results [81]. Diffusion models have emerged as a powerful approach, using noise removal to reconstruct missing regions of images, with models like GLIDE, DALL-E [57] [49] and Stable Diffusion [59] incorporating additional guidance, such as text, for improved control.
@@ 2.2 Inpainting Detection
Detecting inpainted regions has grown crucial with inpainting advancements. Early approaches used patch comparison and connectivity analysis to identify inconsistencies [79, 12, 37, 85]. Deep learning, particularly CNNs, improved detection through high-pass filtering, segmentation networks, and encoder-decoder architectures [91, 35, 40, 32, 63, 39]. Hybrid models combining transformers with CNNs enhanced long-range dependency and texture analysis [92, 89, 76]. Recent methods integrate frequency-domain, compression artifacts, noise, and semantic features [34, 71, 20, 27]. While effective for traditional inpainting and forgeries like copy-move and splicing, few have been tested on AI-based inpainting, except for evaluations on GLIDE-generated forgeries [20, 71, 27], leaving a gap in assessing performance on modern generative techniques.
@@ 2.4 Image Quality and Aesthetic Assessment
Image quality assessment (IQA) evaluates how distortions affect human perception, evolving from handcrafted features [75, 48] to deep learning methods like NIMA [67] and HyperIQA [64]. Recent approaches, such as MUSIQ [29] that utilizes transformers and methods that utilize CLIP [55, 88, 74], leverage multi-scale inputs and vision-language correspondence but often underperform purely visual methods due to reliance on visual-text similarity. Image aesthetic assessment (IAA) is more complex, focusing on attributes like composition and emotional impact (Murray et al., 2012). Deep learning dominates IAA, with methods like NIMA [67] and VILA [30] achieving state-of-the-art results. QALIGN [77] advances both IQA and IAA by leveraging vision-language models to improve generalization without extensive fine-tuning. However, most approaches are task-specific and struggle with generalization, and none of them focus on realism assessment.
@@ 3 Semantically Aligned and Uncertainty Guided AI Image Inpainting
Recent studies have shown that using detailed prompts in AI image generation leads to better inpainting results [60, 47], with human users typically employing complex and carefully crafted prompts [42], while intuitively evaluating image semantics [52] and using a variety of models and generation attempts [68] until they achieve the desired visual result. Inspired by these observations, we argue that automatically generating high-quality inpainted images, deceiving to the human eye, requires addressing two key challenges, i.e. ensuring semantic coherence in manipulations and assessing the realism of the generated images. To this end, we propose SAGI, a model-agnostic pipeline for a) sampling prompts from a distribution that aligns with the human perception of reality and b) assessing at test-time and discarding generated samples that deviate from such a distribution - techniques notably absent in previous works. We consider pre-trained large language models and vison-language models to be some of the best learned approximations currently available for such a distribution, due to the scale of their training data. Ultimately, we employ them in our proposed automated image generation pipeline that consists of two main components: 1) Semantically Aligned Object Replacement that ensures contextually appropriate manipulations and 2) Uncertainty-Guided Deceptiveness Assessment that evaluates the realism of generated images.
(a) ”a delicious apple to add a pop of color to the scene”
(b) ”lush green vines and foliage, intertwining around ancient stone pillars”
(c) ”a fluffy golden retriever”
(d) ”a fluffy white bunny sitting inside a laptop.”
@@
(f) ”a curious otter exploring the grassy field”
Figure 3: Original (with semi-transparent red inpainting mask) and inpainted images from three datasets, with prompts shown below each pair for text-guided models. The first row shows images classified as deceiving by UGDA 3.2, and the second row shows undeceiving images. Each column corresponds to a different dataset: COCO (first), RAISE (second), and OpenImages (third).
@@ 3.1 Semantically Aligned Object Replacement
Generating realistic inpainted imagery requires selecting appropriate objects for replacement and creating contextually relevant prompts that maintain the semantic consistency of the image. To address this challenge, we propose Semantically Aligned Object Replacement (SAOR), a method that automates object selection and prompt generation. Formally, let $I$ be an image and $S(I)$ represent its semantic content. We aim to sample a prominent prompt $p$ and object to be replaced $o$ under the distribution $\mathcal{P}(p,o\mid S(I))$ that characterizes how humans would select regions and generate prompts given the semantics of the image. We approximate $S(I)$ through two components: the image caption $C(I)$ and the set of displayed objects $O(I)$. This decomposition allows us to capture both the local semantics (through the objects) and the global relationships among them (through the caption), providing a convenient, yet descriptive, representation of the image semantics.
@@
Object masks $M(I)=\{m_{1},m_{2},\ldots,m_{n}\}$ and their respective labels $O(I)=\{o_{1},o_{2},\ldots,o_{n}\}$ are obtained directly from labeled data if available; otherwise, a segmentation model $\Phi_{\text{seg}}:I\rightarrow(O(I),M(I))$ is used. Similarly, descriptive captions $C(I)$ are taken directly from labeled data if available; otherwise, they are generated by a captioning model $\Psi_{\text{cap}}$. To approximate $\mathcal{P}$, we leverage large language models as they have been extensively trained on human-generated text, making them ideal candidates to model human reasoning about visual semantics. Specifically:
@@
where $\Theta_{\text{llm}}$ represents the LLM’s conditional probability distribution over generated text. The language model, leveraging the caption to understand the image content, is instructed to produce complex, contextually appropriate prompts, maintaining semantic consistency with the surrounding content. The generated prompt $p$, the mask $m$ corresponding to the object selected by the LLM, and the original image $I$ are then provided to an inpainting model $\Gamma$ to generate the inpainted image $\hat{I}=\Gamma(I,m,p)$. SAOR pipeline is shown in Fig. 1
@@ 3.2 Uncertainty-Guided Deceptiveness Assessment
Assessing the realism of inpainted images at test-time is crucial for discarding unrealistic samples and iteratively refining the generation process. However, quantifying realism is challenging due to its subjective nature and depends on human perception [69]. To approximate this subjective judgment, we propose Uncertainty-Guided Deceptiveness Assessment (UGDA), a process that leverages a large-scale pretrained vision-language model (VLM) to compare inpainted images with their originals. VLMs have also proven to be successful in synthetic image detection [80].
Let $I$ denote the original image and $\hat{I}$ the inpainted version of $I$. Using a VLM $\Omega_{\text{vlm}}$, we perform the assessment in two stages. Based on empirical observations, unrealistic manipulations are often easily identifiable, while realistic ones require a more thorough evaluation. In the first stage of UGDA, $\hat{I}$ is assessed by the VLM for realism. Images that pass this initial realism check undergo a second, more rigorous evaluation.
The second stage leverages the following key idea. The more realistic an inpainted image is, the closer it aligns with the distribution of semantically correct images, which a large-scale pretrained VLM represents. Therefore, the aleatoric uncertainty in a proxy task of ranking the original and its corresponding inpainted image according to their semantical correctness, conditioned on the learned distribution of the VLM, should rise. Consequently, the bigger the aleatoric uncertainty in this proxy task is, the closer the inpainted image is expected to align with the distribution of semantically correct images. To capture this uncertainty increase, we assume that a confident assessment of realism, i.e. with small uncertainty, generated by the VLM, should remain stable under semantically invariant perturbations. Therefore, if the assessment generated by the VLM varies under such perturbations, it indicates uncertainty about the inpainted image’s realism. We argue that such a suitable perturbation is input reordering. Specifically, we perform two assessments with reversed image order:
@@
where $s_{1},s_{2}\in\{c_{I},c_{\hat{I}},c_{\text{both}}\}$ represents the image assessed by the VLM as more realistic, with $c_{I}$ indicating greater realism on the original image, $c_{\hat{I}}$ on the inpainted image, while $c_{\text{both}}$ highlighting that both are equally realistic. Any variation in these two responses or consistently choosing the inpainted image indicates model uncertainty about $\hat{I}$’s realism. The final classification follows these rules: $\hat{I}$ is labeled as deceiving if $(s_{1}=c_{\hat{I}}\vee s_{2}=c_{\hat{I}})\vee(s_{1}=s_{2}=c_{\text{both}})$. Otherwise, $\hat{I}$ is labeled as undeceiving. This approach enhances reliability in distinguishing between deceiving and easily identifiable synthetic images through uncertainty-aware, order-based evaluations. Fig. 3 shows representative examples from UGDA across different sources. The first row demonstrates high-quality inpainting examples classified as deceiving by UGDA, where the manipulations are seamlessly integrated with the original content (see supplementary for detailed classification process). The UGDA pipeline is shown in Fig. 1

### v3

@@ Abstract
Recent advancements in generative AI have made text-guided image inpainting—adding, removing, or altering image regions using textual prompts—widely accessible. However, generating semantically correct photorealistic imagery, typically requires carefully-crafted prompts and iterative refinement by evaluating the realism of the generated content - tasks commonly performed by humans. To automate the generative process, we propose Semantically Aligned and Uncertainty Guided AI Image Inpainting (SAGI), a model-agnostic pipeline, to sample prompts from a distribution that closely aligns with human perception and to evaluate the generated content and discard instances that deviate from such a distribution, which we approximate using pretrained large language models and vision-language models. By applying this pipeline on multiple state-of-the-art inpainting models, we create the SAGI Dataset (SAGI-D), currently the largest and most diverse dataset of AI-generated inpaintings, comprising over 95k inpainted images and a human-evaluated subset. Our experiments show that semantic alignment significantly improves image quality and aesthetics, while uncertainty guidance effectively identifies realistic manipulations — human ability to distinguish inpainted images from real ones drops from 74% to 35% in terms of accuracy, after applying our pipeline. Moreover, using SAGI-D for training several image forensic approaches increases in-domain detection performance on average by 37.4% and out-of-domain generalization by 26.1% in terms of IoU, also demonstrating its utility in countering malicious exploitation of generative AI. Code and dataset are available at https://mever-team.github.io/SAGI/
@@ 1 Introduction
Figure 1: The first row illustrates SAOR: an input image is processed to identify objects (via labels or segmentation) and generate a caption. The language model selects an object and generates a prompt, which, along with the image and mask, guides the inpainting model to produce the result. The second row depicts UGDA: if the inpainted image passes an initial realism check, it undergoes a second evaluation. The VLM compares the inpainted and original images twice, in reverse order. If the VLM’s responses differ (indicating uncertainty) or both favor the inpainted image, it is labeled as deceiving; otherwise, non-deceiving.
@@
Image inpainting—reconstructing image regions—became highly accessible with powerful generative AI tools, enabling non-experts to create photorealistic edits [88]. Text-guided inpainting, that adds, removes, or alters regions using textual prompts, has advanced significantly through models like Stable Diffusion [61], DALL-E [59], and Imagen [64]. Achieving high-quality results often requires several attempts and iterative refinement [44], as advanced models struggle with complex prompts [13]. However, successful generations can be so realistic that humans struggle to distinguish them from real photographs [13].
@@
Automated inpainting generation faces two key challenges. First, while state-of-the-art models excel at manipulation, they require detailed prompts for realistic results—basic object labels often lack sufficient context [43]. Second, automating realism assessment is difficult, as standard quality metrics may not reflect perceived realism. To address these, we propose SAGI (Semantically Aligned and Uncertainty Guided AI Image Inpainting), a unified framework integrating Semantically Aligned Object Replacement (SAOR) and Uncertainty Guided Deceptiveness Assessment (UGDA). SAOR leverages image semantics and large language models (LLMs) to create context-aware prompts, producing higher aesthetic quality than simple object labels and captions. UGDA is a realism assessment method that uses vision-language models (VLMs) to compare inpainted images with originals, identifying convincing manipulations. We validate this approach through a user study, showing it aligns with human perception of image realism. An overview of the approach is presented in Fig. 1.
@@
Using this framework along with multiple state-of-the-art inpainting models, we create the SAGI-D dataset, the largest and most diverse collection of AI-generated inpaintings to date, establishing a new benchmark for inpainting detection research. The dataset contains 95,839 inpainted images from 78,684 originals across three datasets: MS-COCO [40], RAISE [22], and OpenImages [35]. We provide each inpainted image with its original version, inpainting mask, and text prompt. For evaluation, we create both in-domain and out-of-domain testing splits. The in-domain split uses the same LLM and source images as the training set, for familiar data assessment. The out-of-domain split uses different source images and a different LLM, testing generalization to new data. Beyond advancing inpainting generation, the resulting dataset provides a valuable resource for training and evaluating forensic detection methods that can address challenges regarding potential malicious misuse [77, 82].
@@
  • We introduce semantically aligned object replacement to generate semantically coherent prompts.
@@
  • We present the SAGI-D dataset, the largest and most diverse dataset of AI-generated inpainted images to date
@@ 2.1 Image Inpainting
Early image inpainting methods used diffusion for image restoration [11, 12, 10, 14] and exemplar-based approaches for object removal [21, 28, 32, 25]. Recent methods are deep learning-based, using CNNs [9], Fourier Convolutions [68], Transformers [76], or Diffusion Models [16]. GAN-based methods are also widely used, to generate coherent images, with techniques such as Context Encoders [56] and dilated convolutions improving results [85]. Diffusion models have emerged as a powerful approach, reconstructing missing regions via noise removal, with models like GLIDE, DALL-E [60] [52] and Stable Diffusion [62] incorporating additional guidance, such as text, for improved control.
@@ 2.2 Inpainting Detection
Inpainting detection became crucial with inpainting advancements. Early approaches used patch comparison and connectivity analysis to identify inconsistencies [83, 15, 39, 89]. Deep learning, particularly CNNs, improved detection via filtering, segmentation networks, and encoder-decoder architectures [95, 37, 42, 34, 66, 41]. Transformer-CNN hybrids enhanced long-range dependency and texture analysis [96, 93, 80]. Recent methods integrate frequency-domain, compression artifacts, noise, and semantic features [36, 75, 24, 31]. While effective for traditional inpainting and forgeries like copy-move and splicing, few have been tested on AI-based inpainting, except for evaluations on GLIDE-generated forgeries [24, 75, 31], leaving a gap in assessing performance on modern generative techniques.
@@ 2.4 Image Quality and Aesthetic Assessment
Image quality assessment (IQA) evaluates how distortions affect human perception, evolving from handcrafted features [79, 50] to deep learning methods like NIMA [70] and HyperIQA [67]. Recent approaches utilize CLIP [58, 92, 78], leverage multi-scale inputs and vision-language correspondence but often underperform purely visual methods due to reliance on visual-text similarity, which prioritizes semantic alignment over visual fidelity. Image aesthetic assessment (IAA) is more complex, focusing on attributes like composition and emotional impact [51]. Deep learning dominates IAA, with methods like NIMA [70] and VILA [33] achieving state-of-the-art results. QALIGN [81] advances both IQA and IAA by leveraging VLMs to improve generalization without extensive fine-tuning. However, most approaches are task-specific and struggle with generalization, and none of them focus on realism assessment.
@@ 3 Semantically Aligned and Uncertainty Guided AI Image Inpainting
Recent studies have shown that detailed prompts in AI image generation lead to better inpainting results [63, 49], with users employing complex, carefully crafted prompts [44], evaluating image semantics [55] and using multiple models and attempts [71] until achieving desired results. Inspired by these observations, we argue that automatically generating high-quality inpainted images, deceiving to the human eye, requires addressing two key challenges: ensuring semantic coherence in manipulations and assessing the realism of generated images. To this end, we propose SAGI, a model-agnostic pipeline for a) sampling prompts from a distribution that aligns with human perception of reality and b) assessing and discarding generated samples that deviate from such distribution—techniques notably absent in previous works. We consider pre-trained LLMs and VLMs to be the best available approximations for such distribution, due to their training data scale. We employ them in our automated pipeline with two main components: 1) Semantically Aligned Object Replacement for contextually appropriate manipulations and 2) Uncertainty-Guided Deceptiveness Assessment for evaluating realism.
(a) “a delicious apple to add a pop of color to the scene”
(b) “lush green vines and foliage, intertwining around ancient stone pillars”
(c) “a fluffy golden retriever”
(d) “a fluffy white bunny sitting inside a laptop.”
@@
(f) “a curious otter exploring the grassy field”
Figure 3: Original (with semi-transparent red inpainting mask) and inpainted images from three datasets, with prompts shown below each pair for text-guided models. The first row shows images classified as deceiving by UGDA 3.2, and the second row shows non-deceiving images. Each column corresponds to a different dataset: MS-COCO [40] (first), RAISE [22] (second), and OpenImages [35] (third).
@@ 3.1 Semantically Aligned Object Replacement
Generating realistic inpainted imagery requires selecting appropriate objects for replacement and creating contextually relevant prompts that maintain semantic consistency. To address this challenge, we propose Semantically Aligned Object Replacement (SAOR), a method that automates object selection and prompt generation. Formally, let $I$ be an image and $S(I)$ represent its semantic content. We aim to sample a prompt $p$ and object $o$ under the distribution $\mathcal{P}(p,o\mid S(I))$ that characterizes how humans would select regions and generate prompts given image semantics. We approximate $S(I)$ through two components: the image caption $C(I)$ and the set of displayed objects $O(I)$. This decomposition captures both local semantics (through objects) and global relationships (through captions), providing a descriptive representation of image semantics.
@@
Object masks $M(I)=\{m_{1},m_{2},\ldots,m_{n}\}$ and their respective labels $O(I)=\{o_{1},o_{2},\ldots,o_{n}\}$ are obtained from labeled data if available; otherwise, a panoptic segmentation model $\Phi_{\text{seg}}:I\rightarrow(O(I),M(I))$ is used. Similarly, captions $C(I)$ are taken directly from labeled data or generated by a captioning model $\Psi_{\text{cap}}$. To approximate $\mathcal{P}$, we leverage LLMs, which have been extensively trained on human text, making them ideal candidates to model human reasoning about visual semantics. Specifically:
@@
where $\Theta_{\text{llm}}$ represents the LLM’s conditional probability distribution over generated text. The language model, leveraging the caption to understand the image content, is instructed to produce complex, contextually appropriate prompts, maintaining semantic consistency with the surrounding content. The generated prompt $p$, the mask $m$ corresponding to the object selected by the LLM, and the original image $I$ are then provided to an inpainting model $\Gamma$ to generate the inpainted image $\hat{I}=\Gamma(I,m,p)$. The SAOR pipeline is presented at the top of Fig. 1
@@ 3.2 Uncertainty-Guided Deceptiveness Assessment
Assessing the realism of inpainted images at test-time is crucial for discarding unrealistic samples and refining the generation process. However, quantifying realism is challenging due to its subjective nature [73]. To approximate this subjective judgment, we propose Uncertainty-Guided Deceptiveness Assessment (UGDA), which leverages pretrained VLMs, which have proven successful in synthetic image detection [84], to compare inpainted images with their originals.
Let $I$ denote the original image and $\hat{I}$ the inpainted version of $I$. Using a VLM $\Omega_{\text{vlm}}$, we perform a two-stage assessment. Based on empirical observations, unrealistic manipulations are often easily identifiable, while realistic ones require a more thorough evaluation. In the first stage, $\hat{I}$ is assessed for realism. Images passing this check undergo a second, more rigorous evaluation. The second stage leverages this key idea: realistic inpainted images align closer with the distribution of semantically coherent images that a pretrained VLM represents. Therefore, aleatoric uncertainty should increase in a proxy task of ranking the original and its corresponding inpainted image according to their semantic alignment with the learned distribution of the VLM: the bigger the aleatoric uncertainty in this proxy task is, the closer the inpainted image is expected to align with the distribution of semantically correct images. To capture this uncertainty increase, we assume that a confident assessment (i.e.low uncertainty), generated by the VLM, should remain stable under semantically invariant perturbations. If the VLM’s assessment varies under such perturbations, it indicates uncertainty about the inpainted image’s realism. We argue that a suitable perturbation is input reordering. Specifically, we perform two assessments with reversed image order:
@@
where $s_{1},s_{2}\in\{c_{I},c_{\hat{I}},c_{\text{both}}\}$ represents the image assessed as more realistic, with $c_{I}$ indicating the original, $c_{\hat{I}}$ the inpainted image, and $c_{\text{both}}$ equal realism. Variation in responses or consistently choosing the inpainted image indicates uncertainty. The classification rule: $\hat{I}$ is deceiving if $(s_{1}=c_{\hat{I}}\vee s_{2}=c_{\hat{I}})\vee(s_{1}=s_{2}=c_{\text{both}})$. Otherwise, it is non-deceiving. This approach enhances reliability in distinguishing between deceiving and easily identifiable synthetic images through uncertainty-aware, order-based evaluations. Fig. 3 shows representative examples from UGDA across different sources. The first row depicts high-quality inpainting examples classified as deceiving by UGDA, where the manipulations are seamlessly integrated with the original content (see suppl. for detailed classification process). The UGDA pipeline is shown in Fig. 1